OpenStudy (anonymous):

find the area of the region x=2y^2 , x=4+y^2

OpenStudy (australopithecus):

you just need to subtract the integrals of the two functions

OpenStudy (australopithecus):

but first you need to figure out where they intersect

OpenStudy (australopithecus):

so set 2y^(2) = 4 + y^2 2y^(2) - y^(2)= 4 y^(2) = 4 therefore they intersect at -2 and 2

OpenStudy (australopithecus):

I mean when y = -2 and 2

OpenStudy (anonymous):

ok i got that part

OpenStudy (anonymous):

so the integral is from -2 to 2

OpenStudy (anonymous):

ok i got it now

OpenStudy (anonymous):

thank you for the hlep

OpenStudy (anonymous):

*help

OpenStudy (australopithecus):

You need to graph to see which is the top function and which is the bottom function so the points of intersection are: x = 2(2)^(2) = 8 so we have the point (-2, 8) and (-2,8) for x = 2y^(2) now find the other two points x = 4+ (2)^(2) = 8 thus intersection for this function is (-2,8) (-2,8) btw I'm writing x as y and y as x now we just need to analyze the functions by taking the derivative or just analyzing them f(x)=2x^2 and g(x)=4+x^2 we can tell these are parables find the intersections and it should be easy to graph them g(x) has y intercept at (0,4) and no x intercept f(x) has y intercept at 0,0 f(x) has x intercept at 0,0 thus we have the drawing